OpenStudy (anonymous):

6 ------ 5- √3 6 divided by 5 minus square root of 3

OpenStudy (anonymous):

multiple it by \[\frac{ 5+\sqrt{3} }{ 5+\sqrt{3} } \]

OpenStudy (anonymous):

then just simplify